OpenStudy (anonymous):

In a rectangle, one side equals x+10 and its opposite side equals 3y-3. The other side equals 2y and its opposite side equals x+7. Find x and y.

OpenStudy (anonymous):

x+10=3y-3 2y=x+7

OpenStudy (anonymous):

what properties does a rectangle have?

OpenStudy (anonymous):

both sides are parallel, and all angles are right angles.

OpenStudy (anonymous):

is that all?

OpenStudy (anonymous):

opposite sides are congruent, opposite angles are congruent, the diagonals of a rectangle are congruent

OpenStudy (anonymous):

so how might congruence help you here?

OpenStudy (anonymous):

the opposite sides

OpenStudy (anonymous):

How would I do the equation?

OpenStudy (anonymous):

i know its x+10=3y-3, 2y=x+7, how would i solve for x and y?

OpenStudy (anonymous):

solve each equation for y

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

solve each for x, x is easier, get x alone for both equations

OpenStudy (anonymous):

did you get it for either equation?

OpenStudy (anonymous):

x=something

OpenStudy (anonymous):

could you set up the equation?

OpenStudy (anonymous):

you have 2 equations and two unknowns. the first one is x+10=3y-3 can you solve that one for x? so x=.....

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

maybe, tell me how you did it and ill tell you if your right or not

OpenStudy (anonymous):

or tell me y

OpenStudy (anonymous):

 after this you set up a system and solve